Austin is 2 years older than Kori. In 9 years, the sum of their ages will be 54. How old is Austin now?

Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39617)
Answer by ikleyn(52781)

.
let  A  be the Austin age.


Then Kori's age is (A-2) years, according to the condition.


In 9 years Austin age will be (A+9) years, while Kori's age will be ((A-2) + 9) = A + 7 years.


The sum of their age in 9 years is 54 (given), producing the equation


    (A+9) + (A + 7) = 54.


Simplify and solve for A

    2A + 16 = 54

    2A      = 54 - 16

    2A      = 38

     A      = 38/2 = 19.


ANSWER.  Austin is 19 years now.

Solved.